# Improving Diffusion Inverse Problem Solving with Decoupled Noise Annealing

### v1

@@ Abstract
Diffusion models have recently achieved success in solving Bayesian inverse problems with learned data priors. Current methods build on top of the diffusion sampling process, where each denoising step makes small modifications to samples from the previous step. However, this process struggles to correct errors from earlier sampling steps, leading to worse performance in complicated nonlinear inverse problems, such as phase retrieval. To address this challenge, we propose a new method called Decoupled Annealing Posterior Sampling (DAPS) that relies on a novel noise annealing process. Specifically, we decouple consecutive steps in a diffusion sampling trajectory, allowing them to vary considerably from one another while ensuring their time-marginals anneal to the true posterior as we reduce noise levels. This approach enables the exploration of a larger solution space, improving the success rate for accurate reconstructions. We demonstrate that DAPS significantly improves sample quality and stability across multiple image restoration tasks, particularly in complicated nonlinear inverse problems. For example, we achieve a PSNR of 30.72dB on the FFHQ 256 dataset for phase retrieval, which is an improvement of 9.12dB compared to existing methods. Our code is available at the GitHub repository DAPS.
@@ 1 Introduction
(a) Phase retrieval
(b) High dynamic range (HDR) reconstruction
(c) Nonlinear deblurring
(d) Inpainting random pixels
Figure 1: Qualitative comparisons for DAPS and LatentDAPS. All inverse problems considered here are nonlinear, except for random inpainting. Samples from our methods, DAPS and LatentDAPS, are closer to the reference images and have superior visual perceptual quality compared to all existing methods.
@@
Empirically, our method demonstrates significantly improved performance on various inverse problems compared to existing approaches, including DPS [3], DDRM [2], PnP-ADMM [26], PSLD [27], ReSample [25], and RED-diff [28]. Our method can be combined with diffusion models in both raw pixel space and learned latent space, which we refer to as DAPS and LatentDAPS, respectively. As shown in Fig. 1, both methods provide superior reconstructions with improved visual perceptual quality across a wide range of nonlinear inverse problems. Our approach exhibits remarkable stability and sampling quality, particularly for challenging nonlinear inverse problems. On the FFHQ 256 dataset, we achieve a 30.72dB PSNR for noisy phase retrieval, which is 9.12dB higher than all existing methods, while on the ImageNet dataset, which includes a broader range of classes, we achieve a 25.78dB PSNR for phase retrieval, surpassing others by 5.24dB. Additionally, DAPS performs well even with a very small number of neural network evaluations (approximately 100), striking a better balance between efficiency and sample quality.
@@ 2.1 Diffusion Models
Diffusion models [17, 18, 19, 1, 20] generate data by reversing a predefined noising process. Let the data distribution be $p({\mathbf{x}}_{0})$. We can define a series of noisy data distributions $p({\mathbf{x}};\sigma)$ by adding Gaussian noise with a standard deviation of $\sigma$ to the data. These form the time-marginals of a stochastic differential equation (SDE) [20], given by
$\mathrm{d}{\mathbf{x}}_{t}=\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t},$ (1)
where $\sigma_{t}$ is the predefined noise schedule with $\sigma_{0}=0$ and $\sigma_{T}=\sigma_{\max}$, $\dot{\sigma}_{t}$ is the time derivative of $\sigma_{t}$, and ${\mathbf{w}}_{t}$ is a standard Wiener process. We use ${\mathbf{x}}_{t}$ interchangeably with ${\mathbf{x}}_{\sigma_{t}}$. For a sufficiently large $\sigma_{\max}$, the distribution $p({\mathbf{x}};\sigma_{\max})$ converges to pure Gaussian noise $\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$.
To sample from data distribution $p({\mathbf{x}}_{0})$, we first draw an initial sample from $\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$, then solve the reverse SDE
$\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (2)
@@
where $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is the time-dependent score function [17, 1]. Here $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is unknown, but can be approximated by training a diffusion model ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})$ such that ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})\approx\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$.
2.2 Bayesian Inverse Problems with Diffusion Priors
Inverse problems aim to recover data from partial, potentially noisy measurements. Formally, solving an inverse problem involves finding the inversion to a forward model that describes the measurement process. In general, a forward model takes the form of
${\mathbf{y}}=\mathcal{A}({\mathbf{x}}_{0})+{\mathbf{n}},$ (3)
@@ 2.2 Bayesian Inverse Problems with Diffusion Priors
where $\mathcal{A}$ is the measurement function, ${\mathbf{x}}_{0}$ represents the original data, ${\mathbf{y}}$ is the observed measurement, and ${\mathbf{n}}$ symbolizes the noise in the measurement process, often modeled as ${\mathbf{n}}\sim\mathcal{N}({\bm{0}},\beta_{{\mathbf{y}}}^{2}{\bm{I}})$. In a Bayesian framework, ${\mathbf{x}}_{0}$ comes from the posterior distribution $p({\mathbf{x}}_{0}\mid{\mathbf{y}})\propto p({\mathbf{x}}_{0})p({\mathbf{y}}\mid{\mathbf{x}}_{0})$. Here $p({\mathbf{x}}_{0})$ is a prior distribution that can be estimated from a given dataset, and $p({\mathbf{y}}\mid{\mathbf{x}}_{0})=\mathcal{N}(\mathcal{A}({\mathbf{x}}_{0}),\beta_{{\mathbf{y}}}^{2}{\bm{I}})$ models the noisy measurement process.
@@
When the prior $p({\mathbf{x}}_{0})$ is modeled by a pre-trained diffusion model, we can modify Eq. 2 to approximately sample from the posterior distribution following Bayes’ rule, i.e.,
$\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}\Big{(}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})+\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})\Big{)}\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (4)
@@
Here, the noisy likelihood $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})$ is generally intractable. Multiple methods have been proposed to estimate the noisy likelihood [8, 29, 22, 23, 30, 31, 32, 6]. One predominant approach is the DPS algorithm [3], which estimates $p({\mathbf{y}}\mid{\mathbf{x}}_{t})\approx p({\mathbf{y}}\mid{\mathbf{x}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}])$. Another line of work [2, 33, 31] solves linear inverse problems by running the reverse diffusion process in the spectral domain via singular value decomposition (SVD). Other approaches bypass direct computation of this likelihood by interleaving optimization [25, 34, 35, 36, 37] or projection [2, 9, 7, 31] steps with normal diffusion sampling steps.
@@ 3 Method
3.1 Posterior Sampling with Decoupled Noise Annealing

### v2

@@ Abstract
Diffusion models have recently achieved success in solving Bayesian inverse problems with learned data priors. Current methods build on top of the diffusion sampling process, where each denoising step makes small modifications to samples from the previous step. However, this process struggles to correct errors from earlier sampling steps, leading to worse performance in complicated nonlinear inverse problems, such as phase retrieval. To address this challenge, we propose a new method called Decoupled Annealing Posterior Sampling (DAPS) that relies on a novel noise annealing process. Specifically, we decouple consecutive steps in a diffusion sampling trajectory, allowing them to vary considerably from one another while ensuring their time-marginals anneal to the true posterior as we reduce noise levels. This approach enables the exploration of a larger solution space, improving the success rate for accurate reconstructions. We demonstrate that DAPS significantly improves sample quality and stability across multiple image restoration tasks, particularly in complicated nonlinear inverse problems. Our code is available at the GitHub repository DAPS.
Figure 1: Overview of Decoupled Annealing Posterior Sampling (DAPS). Our method provides a flexible and effective framework for solving inverse problems through a decoupled posterior sampling process. In (a)(b), we present DAPS visual results on FFHQ and ImageNet at a resolution of 256, and in (c), on natural images at a resolution of 768. In (d), we display DAPS results on compressed sensing multi-coil MRI (CS-MRI). DAPS effectively addresses nonlinear inverse problems as well as medical imaging MRI challenges. Additionally, DAPS can be enhanced using large-scale latent diffusion models (LDMs) [40], as shown in (c).
@@ 1 Introduction
Empirically, our method demonstrates significantly improved performance on various inverse problems compared to existing approaches, ranging from image restoration to compressed sensing MRI. Our method can be combined with diffusion models in both raw pixel space and learned latent space, which we refer to as DAPS and LatentDAPS, respectively. As shown in Fig. 1, both methods provide superior reconstructions with improved visual perceptual quality across a wide range of nonlinear inverse problems. Our approach exhibits remarkable stability and sampling quality, particularly for challenging nonlinear inverse problems. On the FFHQ 256 dataset, we achieve PSNR values of 30.72 dB and 27.12 dB for the phase retrieval and high dynamic range tasks, representing improvements of 1.98 dB and 4.39 dB over the previous state-of-the-art results, respectively. We apply DAPS to large-scale text-conditioned latent diffusion models, showing its scalability to high-resolution image restoration tasks. Beyond image restoration tasks, DAPS achieves a PSNR of 31.49 dB, showing a 2.70 dB improvement over previous diffusion-based methods. We also demonstrate how DAPS can be applied to inverse problems on categorical data with pre-trained discrete diffusion models. Additionally, DAPS performs well with a small number of neural network evaluations (approximately 100), striking a better balance between efficiency and sample quality.
@@ 2.1 Diffusion Models
Diffusion models [47, 48, 21, 49, 26] generate data by reversing a predefined noising process. Let the data distribution be $p({\mathbf{x}}_{0})$. We can define a series of noisy data distributions $p({\mathbf{x}};\sigma)$ by adding Gaussian noise with a standard deviation of $\sigma$ to the data. These form the time-marginals of a stochastic differential equation (SDE) [26], given by $\mathrm{d}{\mathbf{x}}_{t}=\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}$, where $\sigma_{t}$ is the predefined noise schedule with $\sigma_{0}=0$ and $\sigma_{T}=\sigma_{\max}$, $\dot{\sigma}_{t}$ is the time derivative of $\sigma_{t}$, and ${\mathbf{w}}_{t}$ is a standard Wiener process. We use ${\mathbf{x}}_{t}$ interchangeably with ${\mathbf{x}}_{\sigma_{t}}$. For a sufficiently large $\sigma_{\max}$, the distribution $p({\mathbf{x}};\sigma_{\max})$ converges to pure Gaussian noise $\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$.
@@
To sample from data distribution $p({\mathbf{x}}_{0})$, we first draw an initial sample from $\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$, then solve the reverse SDE: $\mathrm{d}{\mathbf{x}}_{t}=-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t},$ where $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is the time-dependent score function [47, 49]. Here $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$ is unknown, but can be approximated by training a diffusion model ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})$ such that ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},\sigma_{t})\approx\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})$.
2.2 Bayesian Inverse Problems with Diffusion
@@
Inverse problems aim to recover data from partial, potentially noisy measurements. Formally, solving an inverse problem involves finding the inversion to a forward model that describes the measurement process. In general, a forward model takes the form of ${\mathbf{y}}=\mathcal{A}({\mathbf{x}}_{0})+{\mathbf{n}},$ where $\mathcal{A}$ is the measurement function, ${\mathbf{x}}_{0}$ represents the original data, ${\mathbf{y}}$ is the observed measurement, and ${\mathbf{n}}$ symbolizes the noise in the measurement process, often modeled as ${\mathbf{n}}\sim\mathcal{N}({\bm{0}},\beta_{{\mathbf{y}}}^{2}{\bm{I}})$. In a Bayesian framework, ${\mathbf{x}}_{0}$ comes from the posterior distribution $p({\mathbf{x}}_{0}\mid{\mathbf{y}})\propto p({\mathbf{x}}_{0})p({\mathbf{y}}\mid{\mathbf{x}}_{0})$. Here $p({\mathbf{x}}_{0})$ is a prior distribution that can be estimated from a given dataset, and $p({\mathbf{y}}\mid{\mathbf{x}}_{0})=\mathcal{N}(\mathcal{A}({\mathbf{x}}_{0}),\beta_{{\mathbf{y}}}^{2}{\bm{I}})$ models the noisy measurement process.
@@
When the prior $p({\mathbf{x}}_{0})$ is modeled by a pre-trained diffusion model, we can modify reverse SDE to approximately sample from the posterior distribution following Bayes’ rule,
$\mathrm{d}{\mathbf{x}}_{t}=$ $-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t};\sigma_{t})\mathrm{d}t$
$-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}{\mathbf{w}}_{t}.$ (1)
@@
Here, the noisy likelihood $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{t})$ is generally intractable. Multiple methods have been proposed to estimate the noisy likelihood [50, 12, 46, 5, 36, 54, 8, 59, 41]. One predominant approach is the DPS algorithm [13], which estimates $p({\mathbf{y}}\mid{\mathbf{x}}_{t})\approx p({\mathbf{y}}\mid{\mathbf{x}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}])$. Another line of work [30, 29, 54, 7, 18] solves linear inverse problems by running the reverse diffusion in the spectral domain via singular value decomposition (SVD). Other methods bypass direct computation of this likelihood by interleaving optimization [45, 24, 9, 2, 59, 31, 56] or projection [30, 10, 23, 54] steps with normal diffusion sampling steps.
@@ 3 Method
3.1 Decoupled Annealing Posterior Sampling